Given the values of ΔH∘rxn, ΔS∘rxn, and T below, determine ΔSuniv and whether they are spontaneous or nonspontaneous:

A. ΔH∘rxn= 80 kJ , ΔSrxn= 143 J/K , T= 290 K

B. ΔH∘rxn= 80 kJ , ΔSrxn= 143 J/K , T= 764 K

C. ΔH∘rxn= 80 kJ , ΔSrxn=− 143 J/K , T= 290 K

D. ΔH∘rxn=− 80 kJ , ΔSrxn= 143 J/K , T= 396 K
Answer #1

Formulae and statment:

ΔSrxn = ΔSsystem.

ΔSsurrounding = -ΔH∘rxn /T

And,

ΔSuniverse = ΔSsystem + ΔSsurrounding.

For spontaneous reaction,

ΔSuniverse > 0 i.e. +ve

===================================================

Based on above formulae:

A) Given, ΔH∘rxn = 80 kJ = 80000 J , ΔSrxn= 143 J/K , T= 290 K

ΔSrxn= ΔSsystem = 143 J/K

ΔSsurrounding = -ΔH∘rxn/T = -80000/290 = -275.9

ΔSuniverse = ΔSsystem + ΔSsurrounding

ΔSuniverse = (143) + (-275.9)

ΔSuniverse = - 132.9 J/K

ΔSuniverse < 0 i.e. –ve

Hence reaction is non-spontaneous.

============================================

B)Given, ΔH∘rxn= 80 kJ = 80000 J , ΔSrxn= 143 J/K , T= 764 K

ΔSrxn= ΔSsystem = 143 J/K

ΔSsurrounding = -ΔH∘rxn/T = -80000/764 = -104.7

ΔSuniverse = ΔSsystem + ΔSsurrounding

ΔSuniverse = (143) + (-104.7)

ΔSuniverse = +38.3 J/K

ΔSuniverse > 0 i.e. +ve

Hence reaction is spontaneous.

=============================================

C) Given, ΔH∘rxn= 80 kJ = 80000 J , ΔSrxn=− 143 J/K , T= 290 K

ΔSrxn= ΔSsystem = -143 J/K

ΔSsurrounding = -ΔH∘rxn/T = -80000/290 = -275.9 J/K

ΔSuniverse = ΔSsystem + ΔSsurrounding

ΔSuniverse = (-143) + (-275.9)

ΔSuniverse = -418.9 J/K

ΔSuniverse < 0 i.e. -ve

Hence reaction is non-spontaneous.

=============================================

Given, ΔH∘rxn=− 80 kJ = -80000 J, ΔSrxn= 143 J/K , T= 396 K

ΔSrxn= ΔSsystem = 143 J/K

ΔSsurrounding = -ΔH∘rxn/T = -(- 80000)/396 = 80000/396 = 202.2 J/K

ΔSuniverse = ΔSsystem + ΔSsurrounding

ΔSuniverse = (143) + (202.2)

ΔSuniverse = +345.2 J/K

ΔSuniverse > 0 i.e. +ve

Hence reaction is spontaneous.

=============================================